Please help !!!
a sum of 100,000 is invested in a bank at a rate of 6 % p.a .
a)assume the amount of money invested for 5 years .
b)detemine the amountthat will be withdrawn if the compunding will be done
i) once a year ii) monthly iii) continously
formula ;
A = p ( 1+r) ^ n

Answer by stanbon(75887)

a sum of 100,000 is invested in a bank at a rate of 6 % p.a .
a)assume the amount of money invested for 5 years .
For simple interest the formula is:
A=P(1+rt)
A=100,000(1+0.06*5)
=100,000(1+0.30)
=$130,000
------------------
b)detemine the amount that will be withdrawn if the compunding will be done
i) once a year ii) monthly iii) continously
formula ;
Formula for compound interest is:
A=P(1+r/n)^(nt)
i)once a year for 5 years
A=100,000(1+0.06/1)^(1*5)
A=100000(1.06)^5
A=$133822.56
----------------
ii) monthly for 5 years
A=100,000(1+0.06/12)^(12*5)
A=100000(1+0.005)^60
A=$134885.02
-----------------
Formula for continuous compounding is:
A=Pe^(rt)
iii) continuously for 5 years
A=100,000e^(0.06*5)
A=100,000(1.349858808
A=$134985.88
----------
Cheers,
Stan H.